OpenStudy (anonymous):

HELP! WILL GIVE MEDAL!! what is the simplest form of each expression? 1/c^-5 A)C^5 B)5^C C)-C^5 D)-5/c

OpenStudy (anonymous):

1/ c^-5 = 1/(1/c^5) = c^5 Letter A

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

are u good at this stuff? because im really bad at it. and if so could you answer a few more?

OpenStudy (anonymous):

@summertime Oh sure :)

OpenStudy (anonymous):

okay thank u so much!

OpenStudy (anonymous):

what is the value of y^-5/x^-3 when y=-4 x=2 A)10/3 B)128 C)-1/128 D)-128

OpenStudy (anonymous):

(1/y^5) / (1/x^3) Substitute the values. (-1/1024)/(1/8)= -8/1024 = -1/128 C

OpenStudy (anonymous):

Just remember that when the term has a negative exponent, you raise it to that exponent but put it in the denominator. For example, 2^2 = 4 but 2^-2 = 1/4

OpenStudy (anonymous):

is the number written in scientific notation? if not explain 5.1*10^8 A) yes, the number is written in scientific notation B) no, the first factor is not a number between 1 and 10 C) no, it is not written as a number times a power of 10

OpenStudy (anonymous):

A

OpenStudy (anonymous):

what is the number written in standered notation 5.71*10^-3 A)-171.3 B)0.0571 C)0.000571 D)0.00571

OpenStudy (anonymous):

D

OpenStudy (anonymous):

You multiply 5.71 by 1/1000.

OpenStudy (anonymous):

what is the expression written using each base only once? (-4)^-6*(-4)^7 A)4^13 B)-4 C)13 D)1

OpenStudy (anonymous):

I don't understand the question... sorry.

OpenStudy (anonymous):

can you check this for me and see if im right? (3*10^6)(8*10^-4) A)1.1*10^3 B)2.4*10^3 C)1.1*10^-23 D)2.4*10^-23 I put B

OpenStudy (anonymous):

@raechelvictoria (3q^2)^4 A)3q^16 B)81q^6 C)3q^8 D)81q^8 I put b but im not to sure about this one.

OpenStudy (anonymous):

Its letter D.

OpenStudy (anonymous):

It becomes (3^4)([q^2)^4]

OpenStudy (anonymous):

can you help me with just 2 more??? please!

OpenStudy (anonymous):

t^8(t^10)^0 A)t^-80 B)t^80 C)t^8 D)t^-8

OpenStudy (anonymous):

Okay sure :) C because (t^10)^0 becomes t^0 which is 1.

OpenStudy (anonymous):

last one! (3xy^3)^2(xy)^6 A)9x^8y^12 B)3x^8y^12 C)2x^3y12 D)9x^8y^9

OpenStudy (anonymous):

@raechelvictoria

OpenStudy (anonymous):

A